OpenStudy (dbzfan836):

This is emergency! Billy has 1 gallon of paint. He is going to pour it into a paint tray that measures 10 inches wide, 14 inches long, and 4 cm deep. (P.S. 1 gallon = 231 in^3, 1 inch = 2.54 cm) Which of the following scenarios will happen? The paint will not fill the tray by 441 cm^3. The paint will not fill the tray by 10.53 in^3. The paint will fill the tray exactly. The paint will overfill the tray by 10.53 in^3.

OpenStudy (dbzfan836):

When I solved this, my answers were off, very off

OpenStudy (tkhunny):

Okay, what was your result and how did you manage it?

OpenStudy (tkhunny):

10 in * 14 in * 4 cm * (1 in/2.54 cm) = ?? in^3

OpenStudy (dbzfan836):

3612.896

OpenStudy (aum):

10 inches wide, 14 inches long, and 4 cm deep. The last dimension is in cm and it has to be converted to inches: 4 cm is 4/2.54 inches. Now multiply: 10 * 14 * 4/2.54 = ______ in^3 That is the total volume of the tray in in^3. What do you get?

OpenStudy (dbzfan836):

oops, that was in cm, not in.

OpenStudy (dbzfan836):

in in., it is 1422.4in^3

OpenStudy (aum):

Not correct. Just use your calculator to multiply three numbers and divide by 2.54

OpenStudy (dbzfan836):

I first converted all of the in. to cm., then multiplied, and converted it back

OpenStudy (aum):

No need to do that. Most of the dimensions in this problem use inches: 10 inches wide, 14 inches long. Also, a gallon of paint is given as 231 in^3. Only one measurement is in cm. So it would be easier to convert that one measurement to inches and calculate the volume of the tray as suggested earlier.

OpenStudy (tkhunny):

10 in * 14 in * 4 cm * (1 in/2.54 cm) = 10*14*4/2.54 in^3 = 220.47244 in^3 How does that compare to 231 in^3?

OpenStudy (dbzfan836):

the answer would be A

OpenStudy (tkhunny):

?? 231 > 220. There is an overflow problem.